# Training Bilingual LMs with Data Constraints in the Targeted Language

### v1

@@ F.3 Individual Eval Dataset Results
Base PT $23.98\pm 1.25$ $39.14\pm 1.00$ $33.26\pm 0.47$ $60.99\pm 1.14$ $66.30\pm 1.50$ $52.01\pm 1.40$ $45.95$
Base EN $25.34\pm 1.27$ $48.02\pm 1.03$ $47.19\pm 0.50$ $71.60\pm 1.05$ $74.00\pm 1.39$ $50.04\pm 1.41$ $52.70$
Base EN + OH/ELI5 Filter $29.44\pm 1.33$ $53.79\pm 1.02$ $49.53\pm 0.50$ $71.49\pm 1.05$ $80.30\pm 1.26$ $55.33\pm 1.40$ $56.65$
ARC EN $30.72\pm 1.35$ $57.83\pm 1.01$ $48.79\pm 0.50$ $73.01\pm 1.04$ $78.30\pm 1.30$ $52.64\pm 1.40$ $56.88$
SciQ+Inst $29.35\pm 1.33$ $57.45\pm 1.01$ $48.26\pm 0.50$ $71.55\pm 1.05$ $78.80\pm 1.29$ $53.99\pm 1.40$ $56.56$
FWE $35.07\pm 1.39$ $64.39\pm 0.98$ $54.81\pm 0.50$ $72.69\pm 1.04$ $82.10\pm 1.21$ $57.62\pm 1.39$ $61.11$
Base PT $30.25\pm 1.36$ $43.55\pm 1.04$ $42.36\pm 0.51$ $64.80\pm 1.11$ $69.88\pm 1.49$ $52.11\pm 1.42$ $50.49$
Base EN $24.06\pm 1.26$ $38.04\pm 1.02$ $36.04\pm 0.50$ $59.68\pm 1.14$ $59.18\pm 1.59$ $51.06\pm 1.42$ $44.68$
Base EN + OH/ELI5 Filter $25.37\pm 1.29$ $39.59\pm 1.03$ $37.15\pm 0.50$ $58.71\pm 1.15$ $66.11\pm 1.53$ $48.70\pm 1.42$ $45.94$
ARC EN $27.90\pm 1.32$ $38.79\pm 1.02$ $36.87\pm 0.50$ $60.07\pm 1.14$ $64.32\pm 1.55$ $49.35\pm 1.42$ $46.22$
SciQ+Inst $27.03\pm 1.31$ $40.38\pm 1.03$ $37.72\pm 0.50$ $60.72\pm 1.14$ $66.00\pm 1.54$ $50.73\pm 1.42$ $47.10$
FWE $28.86\pm 1.34$ $43.11\pm 1.04$ $39.81\pm 0.51$ $60.34\pm 1.14$ $66.84\pm 1.53$ $50.57\pm 1.42$ $48.25$
@@
Base IT $21.59\pm 1.20$ $35.77\pm 0.98$ $32.54\pm 0.47$ $59.09\pm 1.15$ $63.70\pm 1.52$ $51.07\pm 1.40$ $43.96$
Base EN $26.11\pm 1.28$ $48.53\pm 1.03$ $47.16\pm 0.50$ $71.71\pm 1.05$ $73.50\pm 1.40$ $53.35\pm 1.40$ $53.39$
Base EN + OH/ELI5 Filter $31.14\pm 1.35$ $57.37\pm 1.01$ $50.89\pm 0.50$ $72.42\pm 1.04$ $79.50\pm 1.28$ $53.12\pm 1.40$ $57.40$
ARC EN $30.46\pm 1.34$ $56.06\pm 1.02$ $48.78\pm 0.50$ $73.01\pm 1.04$ $75.90\pm 1.35$ $52.57\pm 1.40$ $56.13$
SciQ+Inst $30.38\pm 1.34$ $56.78\pm 1.02$ $48.49\pm 0.50$ $71.27\pm 1.06$ $80.20\pm 1.26$ $54.85\pm 1.40$ $56.99$
FWE $37.03\pm 1.41$ $65.61\pm 0.97$ $54.91\pm 0.50$ $72.74\pm 1.04$ $84.20\pm 1.15$ $54.85\pm 1.40$ $61.56$
Base IT $26.33\pm 1.30$ $40.38\pm 1.03$ $39.89\pm 0.51$ $64.74\pm 1.11$ $61.76\pm 1.58$ $51.42\pm 1.42$ $47.42$
Base EN $25.81\pm 1.29$ $36.02\pm 1.01$ $35.29\pm 0.50$ $58.65\pm 1.15$ $58.82\pm 1.60$ $52.23\pm 1.42$ $44.47$
Base EN + OH/ELI5 Filter $24.93\pm 1.28$ $32.01\pm 0.98$ $31.76\pm 0.49$ $54.95\pm 1.16$ $61.97\pm 1.57$ $53.20\pm 1.42$ $43.14$
ARC EN $25.72\pm 1.29$ $39.15\pm 1.02$ $36.18\pm 0.50$ $60.55\pm 1.14$ $63.24\pm 1.56$ $51.34\pm 1.42$ $46.03$
SciQ+Inst $26.07\pm 1.30$ $40.03\pm 1.03$ $36.67\pm 0.50$ $58.76\pm 1.15$ $62.29\pm 1.57$ $50.93\pm 1.42$ $45.79$
FWE $29.90\pm 1.35$ $41.04\pm 1.03$ $38.24\pm 0.51$ $62.68\pm 1.13$ $61.03\pm 1.58$ $51.18\pm 1.42$ $47.34$
@@
Base KO $22.10\pm 1.21$ $37.46\pm 0.99$ $28.99\pm 0.45$ $59.19\pm 1.15$ $60.70\pm 1.55$ $51.62\pm 1.40$ $43.34$
Base EN $26.02\pm 1.28$ $47.47\pm 1.02$ $47.71\pm 0.50$ $71.11\pm 1.06$ $75.00\pm 1.37$ $52.72\pm 1.40$ $53.34$
Base EN + OH/ELI5 Filter $30.03\pm 1.34$ $56.69\pm 1.02$ $50.38\pm 0.50$ $71.76\pm 1.05$ $78.90\pm 1.29$ $53.20\pm 1.40$ $56.83$
ARC EN $30.03\pm 1.34$ $57.37\pm 1.01$ $49.41\pm 0.50$ $73.78\pm 1.03$ $78.20\pm 1.31$ $53.12\pm 1.40$ $56.98$
SciQ+Inst $30.97\pm 1.35$ $58.54\pm 1.01$ $48.16\pm 0.50$ $71.49\pm 1.05$ $80.30\pm 1.26$ $52.33\pm 1.40$ $56.97$
FWE $36.01\pm 1.40$ $63.59\pm 0.99$ $54.41\pm 0.50$ $72.96\pm 1.04$ $80.10\pm 1.26$ $55.56\pm 1.40$ $60.44$
Base KO $28.07\pm 1.33$ $42.18\pm 1.04$ $35.48\pm 0.48$ $60.66\pm 1.14$ $71.64\pm 1.47$ $49.28\pm 1.46$ $47.89$
Base EN $22.76\pm 1.24$ $33.82\pm 0.99$ $28.88\pm 0.45$ $55.22\pm 1.16$ $56.08\pm 1.62$ $51.23\pm 1.46$ $41.33$
Base EN + OH/ELI5 Filter $24.85\pm 1.28$ $34.13\pm 1.00$ $29.35\pm 0.45$ $55.55\pm 1.16$ $63.92\pm 1.56$ $52.34\pm 1.46$ $43.35$
ARC EN $25.28\pm 1.28$ $33.73\pm 0.99$ $29.33\pm 0.45$ $53.86\pm 1.16$ $64.44\pm 1.56$ $51.15\pm 1.46$ $42.97$
SciQ+Inst $23.19\pm 1.25$ $34.26\pm 1.00$ $30.01\pm 0.46$ $55.60\pm 1.16$ $64.02\pm 1.56$ $50.04\pm 1.46$ $42.86$
FWE $22.93\pm 1.24$ $32.32\pm 0.98$ $29.09\pm 0.45$ $53.26\pm 1.16$ $64.76\pm 1.55$ $48.43\pm 1.46$ $41.80$
@@
Base JA $23.46\pm 1.24$ $37.42\pm 0.99$ $29.07\pm 0.45$ $59.14\pm 1.15$ $68.10\pm 1.47$ $50.28\pm 1.41$ $44.58$
Base EN $26.19\pm 1.28$ $49.24\pm 1.03$ $48.69\pm 0.50$ $72.63\pm 1.04$ $74.40\pm 1.38$ $53.28\pm 1.40$ $54.07$
Base EN + OH/ELI5 Filter $29.78\pm 1.34$ $55.30\pm 1.02$ $50.26\pm 0.50$ $72.14\pm 1.05$ $78.90\pm 1.29$ $52.72\pm 1.40$ $56.52$
ARC EN $29.35\pm 1.33$ $56.94\pm 1.02$ $49.05\pm 0.50$ $73.78\pm 1.03$ $79.80\pm 1.27$ $54.22\pm 1.40$ $57.19$
SciQ+Inst $30.38\pm 1.34$ $56.14\pm 1.02$ $48.32\pm 0.50$ $72.42\pm 1.04$ $78.60\pm 1.30$ $52.64\pm 1.40$ $56.42$
FineWeb EDU $34.73\pm 1.39$ $63.59\pm 0.99$ $54.80\pm 0.50$ $72.91\pm 1.04$ $82.20\pm 1.21$ $56.43\pm 1.39$ $60.78$
Base JA $25.28\pm 1.28$ $40.69\pm 1.03$ $35.11\pm 0.48$ $58.98\pm 1.15$ $69.98\pm 1.51$ $51.19\pm 1.51$ $46.87$
Base EN $25.28\pm 1.28$ $36.15\pm 1.01$ $31.71\pm 0.46$ $56.26\pm 1.16$ $66.41\pm 1.55$ $50.82\pm 1.51$ $44.44$
Base EN + OH/ELI5 Filter $26.50\pm 1.30$ $36.99\pm 1.01$ $31.11\pm 0.46$ $56.96\pm 1.16$ $67.60\pm 1.54$ $50.18\pm 1.51$ $44.89$
ARC EN $27.55\pm 1.32$ $37.43\pm 1.02$ $31.56\pm 0.46$ $57.29\pm 1.15$ $65.55\pm 1.56$ $48.72\pm 1.51$ $44.68$
SciQ+Inst $27.38\pm 1.32$ $36.42\pm 1.01$ $31.44\pm 0.46$ $56.58\pm 1.16$ $67.28\pm 1.54$ $50.36\pm 1.51$ $44.91$
FineWeb EDU $25.28\pm 1.28$ $33.73\pm 0.99$ $30.91\pm 0.46$ $56.09\pm 1.16$ $63.28\pm 1.58$ $47.08\pm 1.51$ $42.73$
@@
Base ZH $21.16\pm 1.19$ $33.16\pm 0.97$ $27.63\pm 0.45$ $55.98\pm 1.16$ $56.60\pm 1.57$ $49.09\pm 1.41$ $40.61$
Base EN $25.85\pm 1.28$ $47.90\pm 1.03$ $48.54\pm 0.50$ $71.93\pm 1.05$ $73.90\pm 1.39$ $52.64\pm 1.40$ $53.46$
Base EN + OH/ELI5 Filter $28.84\pm 1.32$ $55.85\pm 1.02$ $49.63\pm 0.50$ $71.22\pm 1.06$ $78.70\pm 1.30$ $53.12\pm 1.40$ $56.23$
ARC EN $30.72\pm 1.35$ $57.49\pm 1.01$ $48.38\pm 0.50$ $73.29\pm 1.03$ $80.10\pm 1.26$ $53.75\pm 1.40$ $57.29$
SciQ+Inst $30.38\pm 1.34$ $56.31\pm 1.02$ $47.66\pm 0.50$ $71.16\pm 1.06$ $78.10\pm 1.31$ $52.49\pm 1.40$ $56.02$
FWE $36.18\pm 1.40$ $67.13\pm 0.96$ $54.07\pm 0.50$ $73.99\pm 1.02$ $80.90\pm 1.24$ $55.96\pm 1.40$ $61.37$
Base ZH $25.65\pm 1.29$ $38.88\pm 1.02$ $33.07\pm 0.49$ $56.37\pm 1.16$ $69.90\pm 1.45$ $49.86\pm 1.54$ $45.62$
Base EN $25.22\pm 1.28$ $36.77\pm 1.01$ $31.62\pm 0.48$ $56.09\pm 1.16$ $68.00\pm 1.48$ $48.35\pm 1.54$ $44.34$
Base EN + OH/ELI5 Filter $23.56\pm 1.25$ $38.22\pm 1.02$ $32.45\pm 0.49$ $54.68\pm 1.16$ $68.50\pm 1.47$ $52.79\pm 1.53$ $45.03$
ARC EN $23.21\pm 1.25$ $37.52\pm 1.02$ $32.06\pm 0.48$ $55.44\pm 1.16$ $70.90\pm 1.44$ $48.54\pm 1.54$ $44.61$
SciQ+Inst $22.16\pm 1.23$ $38.00\pm 1.02$ $32.24\pm 0.49$ $54.03\pm 1.16$ $68.60\pm 1.47$ $53.07\pm 1.53$ $44.68$
FWE $25.04\pm 1.28$ $36.50\pm 1.01$ $31.89\pm 0.48$ $54.62\pm 1.16$ $66.30\pm 1.50$ $51.94\pm 1.54$ $44.38$
@@
Appendix G Results for Individual datasets for German Target Language Models

### v2

@@ H.5 Individual Eval Dataset Results
PT (S) $23.38\pm 1.24$ $26.77\pm 0.91$ $26.01\pm 0.44$ $53.48\pm 1.16$ $33.50\pm 1.49$ $50.91\pm 1.41$ $35.67$
PT (L) $23.98\pm 1.25$ $39.14\pm 1.00$ $33.26\pm 0.47$ $60.99\pm 1.14$ $66.30\pm 1.50$ $52.01\pm 1.40$ $45.95$
PT (S) EN(L) $25.34\pm 1.27$ $48.02\pm 1.03$ $47.19\pm 0.50$ $71.60\pm 1.05$ $74.00\pm 1.39$ $50.04\pm 1.41$ $52.70$
PT (S) DCLM EN (L) $29.44\pm 1.33$ $53.79\pm 1.02$ $49.53\pm 0.50$ $71.49\pm 1.05$ $80.30\pm 1.26$ $55.33\pm 1.40$ $56.65$
PT (S) ARC EN (L) $30.72\pm 1.35$ $57.83\pm 1.01$ $48.79\pm 0.50$ $73.01\pm 1.04$ $78.30\pm 1.30$ $52.64\pm 1.40$ $56.88$
PT (S) S+I EN(L) $29.35\pm 1.33$ $57.45\pm 1.01$ $48.26\pm 0.50$ $71.55\pm 1.05$ $78.80\pm 1.29$ $53.99\pm 1.40$ $56.56$
PT (S) FWE (L) $35.07\pm 1.39$ $64.39\pm 0.98$ $54.81\pm 0.50$ $72.69\pm 1.04$ $82.10\pm 1.21$ $57.62\pm 1.39$ $61.11$
PT (S) $25.37\pm 1.29$ $30.74\pm 0.97$ $29.18\pm 0.47$ $55.39\pm 1.16$ $41.97\pm 1.60$ $50.00\pm 1.43$ $38.77$
PT (L) $30.25\pm 1.36$ $43.55\pm 1.04$ $42.36\pm 0.51$ $64.80\pm 1.11$ $69.88\pm 1.49$ $52.11\pm 1.42$ $50.49$
PT (S) EN(L) $24.06\pm 1.26$ $38.04\pm 1.02$ $36.04\pm 0.50$ $59.68\pm 1.14$ $59.18\pm 1.59$ $51.06\pm 1.42$ $44.68$
PT (S) DCLM EN (L) $25.37\pm 1.29$ $39.59\pm 1.03$ $37.15\pm 0.50$ $58.71\pm 1.15$ $66.11\pm 1.53$ $48.70\pm 1.42$ $45.94$
PT (S) ARC EN (L) $27.90\pm 1.32$ $38.79\pm 1.02$ $36.87\pm 0.50$ $60.07\pm 1.14$ $64.32\pm 1.55$ $49.35\pm 1.42$ $46.22$
PT (S) S+I EN(L) $27.03\pm 1.31$ $40.38\pm 1.03$ $37.72\pm 0.50$ $60.72\pm 1.14$ $66.00\pm 1.54$ $50.73\pm 1.42$ $47.10$
PT (S) FWE (L) $28.86\pm 1.34$ $43.11\pm 1.04$ $39.81\pm 0.51$ $60.34\pm 1.14$ $66.84\pm 1.53$ $50.57\pm 1.42$ $48.25$
@@
IT (S) $23.46\pm 1.24$ $27.23\pm 0.91$ $26.30\pm 0.44$ $53.21\pm 1.16$ $32.40\pm 1.48$ $49.88\pm 1.41$ $35.41$
IT (L) $21.59\pm 1.20$ $35.77\pm 0.98$ $32.54\pm 0.47$ $59.09\pm 1.15$ $63.70\pm 1.52$ $51.07\pm 1.40$ $43.96$
IT (S) EN(L) $26.11\pm 1.28$ $48.53\pm 1.03$ $47.16\pm 0.50$ $71.71\pm 1.05$ $73.50\pm 1.40$ $53.35\pm 1.40$ $53.39$
IT (S) DCLM EN (L) $31.14\pm 1.35$ $57.37\pm 1.01$ $50.89\pm 0.50$ $72.42\pm 1.04$ $79.50\pm 1.28$ $53.12\pm 1.40$ $57.40$
IT (S) ARC EN (L) $30.46\pm 1.34$ $56.06\pm 1.02$ $48.78\pm 0.50$ $73.01\pm 1.04$ $75.90\pm 1.35$ $52.57\pm 1.40$ $56.13$
IT (S) S+I EN(L) $30.38\pm 1.34$ $56.78\pm 1.02$ $48.49\pm 0.50$ $71.27\pm 1.06$ $80.20\pm 1.26$ $54.85\pm 1.40$ $56.99$
IT (S) FWE (L) $37.03\pm 1.41$ $65.61\pm 0.97$ $54.91\pm 0.50$ $72.74\pm 1.04$ $84.20\pm 1.15$ $54.85\pm 1.40$ $61.56$
IT (S) $23.63\pm 1.25$ $29.19\pm 0.95$ $28.49\pm 0.47$ $56.31\pm 1.16$ $36.76\pm 1.56$ $48.01\pm 1.42$ $37.07$
IT (L) $26.33\pm 1.30$ $40.38\pm 1.03$ $39.89\pm 0.51$ $64.74\pm 1.11$ $61.76\pm 1.58$ $51.42\pm 1.42$ $47.42$
IT (S) EN(L) $25.81\pm 1.29$ $36.02\pm 1.01$ $35.29\pm 0.50$ $58.65\pm 1.15$ $58.82\pm 1.60$ $52.23\pm 1.42$ $44.47$
IT (S) DCLM EN (L) $24.93\pm 1.28$ $32.01\pm 0.98$ $31.76\pm 0.49$ $54.95\pm 1.16$ $61.97\pm 1.57$ $53.20\pm 1.42$ $43.14$
IT (S) ARC EN (L) $25.72\pm 1.29$ $39.15\pm 1.02$ $36.18\pm 0.50$ $60.55\pm 1.14$ $63.24\pm 1.56$ $51.34\pm 1.42$ $46.03$
IT (S) S+I EN(L) $26.07\pm 1.30$ $40.03\pm 1.03$ $36.67\pm 0.50$ $58.76\pm 1.15$ $62.29\pm 1.57$ $50.93\pm 1.42$ $45.79$
IT (S) FWE (L) $29.90\pm 1.35$ $41.04\pm 1.03$ $38.24\pm 0.51$ $62.68\pm 1.13$ $61.03\pm 1.58$ $51.18\pm 1.42$ $47.34$
@@
KO (S) $23.04\pm 1.23$ $26.14\pm 0.90$ $25.41\pm 0.43$ $50.98\pm 1.17$ $29.50\pm 1.44$ $52.57\pm 1.40$ $34.61$
KO (L) $22.10\pm 1.21$ $37.46\pm 0.99$ $28.99\pm 0.45$ $59.19\pm 1.15$ $60.70\pm 1.55$ $51.62\pm 1.40$ $43.34$
KO (S) EN(L) $26.02\pm 1.28$ $47.47\pm 1.02$ $47.71\pm 0.50$ $71.11\pm 1.06$ $75.00\pm 1.37$ $52.72\pm 1.40$ $53.34$
KO (S) DCLM EN (L) $30.03\pm 1.34$ $56.69\pm 1.02$ $50.38\pm 0.50$ $71.76\pm 1.05$ $78.90\pm 1.29$ $53.20\pm 1.40$ $56.83$
KO (S) ARC EN (L) $30.03\pm 1.34$ $57.37\pm 1.01$ $49.41\pm 0.50$ $73.78\pm 1.03$ $78.20\pm 1.31$ $53.12\pm 1.40$ $56.98$
KO (S) S+I EN(L) $30.97\pm 1.35$ $58.54\pm 1.01$ $48.16\pm 0.50$ $71.49\pm 1.05$ $80.30\pm 1.26$ $52.33\pm 1.40$ $56.97$
KO (S) FWE (L) $36.01\pm 1.40$ $63.59\pm 0.99$ $54.41\pm 0.50$ $72.96\pm 1.04$ $80.10\pm 1.26$ $55.56\pm 1.40$ $60.44$
KO (S) $24.41\pm 1.27$ $28.84\pm 0.95$ $27.63\pm 0.45$ $52.39\pm 1.17$ $47.41\pm 1.63$ $50.89\pm 1.46$ $38.60$
KO (L) $28.07\pm 1.33$ $42.18\pm 1.04$ $35.48\pm 0.48$ $60.66\pm 1.14$ $71.64\pm 1.47$ $49.28\pm 1.46$ $47.89$
KO (S) EN(L) $22.76\pm 1.24$ $33.82\pm 0.99$ $28.88\pm 0.45$ $55.22\pm 1.16$ $56.08\pm 1.62$ $51.23\pm 1.46$ $41.33$
KO (S) DCLM EN (L) $24.85\pm 1.28$ $34.13\pm 1.00$ $29.35\pm 0.45$ $55.55\pm 1.16$ $63.92\pm 1.56$ $52.34\pm 1.46$ $43.35$
KO (S) ARC EN (L) $25.28\pm 1.28$ $33.73\pm 0.99$ $29.33\pm 0.45$ $53.86\pm 1.16$ $64.44\pm 1.56$ $51.15\pm 1.46$ $42.97$
KO (S) S+I EN(L) $23.19\pm 1.25$ $34.26\pm 1.00$ $30.01\pm 0.46$ $55.60\pm 1.16$ $64.02\pm 1.56$ $50.04\pm 1.46$ $42.86$
KO (S) FWE (L) $22.93\pm 1.24$ $32.32\pm 0.98$ $29.09\pm 0.45$ $53.26\pm 1.16$ $64.76\pm 1.55$ $48.43\pm 1.46$ $41.80$
@@
JA (S) $22.61\pm 1.22$ $28.49\pm 0.93$ $26.29\pm 0.44$ $53.21\pm 1.16$ $33.80\pm 1.50$ $51.46\pm 1.40$ $35.98$
JA (L) $23.46\pm 1.24$ $37.42\pm 0.99$ $29.07\pm 0.45$ $59.14\pm 1.15$ $68.10\pm 1.47$ $50.28\pm 1.41$ $44.58$
JA (S) EN(L) $26.19\pm 1.28$ $49.24\pm 1.03$ $48.69\pm 0.50$ $72.63\pm 1.04$ $74.40\pm 1.38$ $53.28\pm 1.40$ $54.07$
JA (S) DCLM EN (L) $29.78\pm 1.34$ $55.30\pm 1.02$ $50.26\pm 0.50$ $72.14\pm 1.05$ $78.90\pm 1.29$ $52.72\pm 1.40$ $56.52$
JA (S) ARC EN (L) $29.35\pm 1.33$ $56.94\pm 1.02$ $49.05\pm 0.50$ $73.78\pm 1.03$ $79.80\pm 1.27$ $54.22\pm 1.40$ $57.19$
JA (S) S+I EN(L) $30.38\pm 1.34$ $56.14\pm 1.02$ $48.32\pm 0.50$ $72.42\pm 1.04$ $78.60\pm 1.30$ $52.64\pm 1.40$ $56.42$
JA (S) FWE EN (L) $36.09\pm 1.40$ $63.93\pm 0.99$ $54.74\pm 0.50$ $73.18\pm 1.03$ $80.00\pm 1.27$ $55.80\pm 1.40$ $60.62$
JA (S) $22.58\pm 1.24$ $30.07\pm 0.96$ $29.32\pm 0.45$ $53.81\pm 1.16$ $41.79\pm 1.62$ $48.81\pm 1.51$ $37.73$
JA (L) $25.28\pm 1.28$ $40.69\pm 1.03$ $35.11\pm 0.48$ $58.98\pm 1.15$ $69.98\pm 1.51$ $51.19\pm 1.51$ $46.87$
JA (S) EN(L) $25.28\pm 1.28$ $36.15\pm 1.01$ $31.71\pm 0.46$ $56.26\pm 1.16$ $66.41\pm 1.55$ $50.82\pm 1.51$ $44.44$
JA (S) DCLM EN (L) $26.50\pm 1.30$ $36.99\pm 1.01$ $31.11\pm 0.46$ $56.96\pm 1.16$ $67.60\pm 1.54$ $50.18\pm 1.51$ $44.89$
JA (S) ARC EN (L) $27.55\pm 1.32$ $37.43\pm 1.02$ $31.56\pm 0.46$ $57.29\pm 1.15$ $65.55\pm 1.56$ $48.72\pm 1.51$ $44.68$
JA (S) S+I EN(L) $27.38\pm 1.32$ $36.42\pm 1.01$ $31.44\pm 0.46$ $56.58\pm 1.16$ $67.28\pm 1.54$ $50.36\pm 1.51$ $44.91$
JA (S) FWE EN (L) $25.89\pm 1.29$ $35.71\pm 1.01$ $31.06\pm 0.46$ $56.20\pm 1.16$ $66.09\pm 1.56$ $49.64\pm 1.51$ $44.10$
@@
ZH (S) $22.10\pm 1.21$ $26.14\pm 0.90$ $25.74\pm 0.44$ $52.50\pm 1.17$ $31.10\pm 1.46$ $50.43\pm 1.41$ $34.67$
ZH (L) $21.16\pm 1.19$ $33.16\pm 0.97$ $27.63\pm 0.45$ $55.98\pm 1.16$ $56.60\pm 1.57$ $49.09\pm 1.41$ $40.61$
ZH (S) EN(L) $25.85\pm 1.28$ $47.90\pm 1.03$ $48.54\pm 0.50$ $71.93\pm 1.05$ $73.90\pm 1.39$ $52.64\pm 1.40$ $53.46$
ZH (S) DCLM EN (L) $28.84\pm 1.32$ $55.85\pm 1.02$ $49.63\pm 0.50$ $71.22\pm 1.06$ $78.70\pm 1.30$ $53.12\pm 1.40$ $56.23$
ZH (S) ARC EN (L) $30.72\pm 1.35$ $57.49\pm 1.01$ $48.38\pm 0.50$ $73.29\pm 1.03$ $80.10\pm 1.26$ $53.75\pm 1.40$ $57.29$
ZH (S) S+I EN(L) $30.38\pm 1.34$ $56.31\pm 1.02$ $47.66\pm 0.50$ $71.16\pm 1.06$ $78.10\pm 1.31$ $52.49\pm 1.40$ $56.02$
ZH (S) FWE (L) $36.18\pm 1.40$ $67.13\pm 0.96$ $54.07\pm 0.50$ $73.99\pm 1.02$ $80.90\pm 1.24$ $55.96\pm 1.40$ $61.37$
ZH (S) $22.77\pm 1.24$ $28.58\pm 0.95$ $28.46\pm 0.47$ $50.71\pm 1.17$ $40.80\pm 1.55$ $48.06\pm 1.54$ $36.56$
ZH (L) $25.65\pm 1.29$ $38.88\pm 1.02$ $33.07\pm 0.49$ $56.37\pm 1.16$ $69.90\pm 1.45$ $49.86\pm 1.54$ $45.62$
ZH (S) EN(L) $25.22\pm 1.28$ $36.77\pm 1.01$ $31.62\pm 0.48$ $56.09\pm 1.16$ $68.00\pm 1.48$ $48.35\pm 1.54$ $44.34$
ZH (S) DCLM EN (L) $23.56\pm 1.25$ $38.22\pm 1.02$ $32.45\pm 0.49$ $54.68\pm 1.16$ $68.50\pm 1.47$ $52.79\pm 1.53$ $45.03$
ZH (S) ARC EN (L) $23.21\pm 1.25$ $37.52\pm 1.02$ $32.06\pm 0.48$ $55.44\pm 1.16$ $70.90\pm 1.44$ $48.54\pm 1.54$ $44.61$
ZH (S) S+I EN(L) $22.16\pm 1.23$ $38.00\pm 1.02$ $32.24\pm 0.49$ $54.03\pm 1.16$ $68.60\pm 1.47$ $53.07\pm 1.53$ $44.68$
ZH (S) FWE (L) $25.04\pm 1.28$ $36.50\pm 1.01$ $31.89\pm 0.48$ $54.62\pm 1.16$ $66.30\pm 1.50$ $51.94\pm 1.54$ $44.38$
@@
Appendix I Results for Individual datasets for German Target Language Models